OpenStudy (anonymous):

What is the solution to the inequality below? x2≥16

OpenStudy (anonymous):

\[x^{2}\ge16 \] \[\sqrt{x^{2}}\ge \sqrt{16}\] \[x \ge \pm4\] So, the solution lies in the interval [-4,4]

OpenStudy (anonymous):

so would it be 'And' or 'Or' ?

OpenStudy (anonymous):

its an "OR"

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

No problem :)